You inherit $11,000, which will be paid to you after the probate hearing. You expect to receive the money one year from today. The interest rate is 10%. What is the present value of your inheritance?
Answer #1

Answer

Formula :

PV = A/(1 + r)n

where, PV = Present Value, A = amount after n years = 11000 , r = interest rate = 10% = 0.10 and n = time period = 1

=> PV = 11000/(1 + 0.10)1= 10,000

Hence, the present value of your inheritance is $10000